# HEIGHT: HEterogeneous Interaction GrapH Transformer for Robot Navigation in Crowded and Constrained Environments

### v1

@@ V Simulation Experiments
V-A Simulation environment
We conduct simulation experiments in random environment in Fig. 5(a) developed with PyBullet [58]. The robot, humans, and static obstacles are in a 12 m×$\times$12 m arena. In each episode, rectangular obstacles are initialized with random shapes and random poses. The width and the length of each static obstacle are sampled from $\mathcal{N}(1,0.6^{2})$ and are limited in $[0.1,5]$ meters. The initial positions of the humans and the robot are also randomized. The starting and goal positions of the robot are randomly sampled inside the arena. The distance between the robot’s starting and the goal positions is between 5 m and 6 m. Some humans are dynamic and some are static. The goals of dynamic humans are set on the opposite side of their initial positions to create circle-crossing scenarios. In training, the number of humans varies from 5 to 9 and the number of static rectangular obstacles varies from 8 to 12. Among all humans, 0-2 of them are static and the rest are dynamic. In testing, the number of humans and obstacles are shown in the first column of Table I. The heights of humans and obstacles are all above the height of the robot LiDAR to ensure detectability.
@@ V-B1 Baselines
We compare the performance of our method with the following baselines:
  • $\textrm{A}^{*}$+ CNN [7] is a hybrid method. With a map of the environment, $\textrm{A}^{*}$ is the global planner and generates 6 waypoints in the beginning of an episode. The inputs to the robot RL policy are a 2D LiDAR point cloud, the robot state, the waypoints, and the robot’s final goal. No human detections are used and human features are included in the point cloud. The point cloud is passed through a CNN local planner, whose output is concatenated with the embedding of other inputs. In addition to the reward function in Eq. 1, the robot receives a small reward of 2 if it arrives at any waypoint in sequence.
In summary, DWA, $\textrm{A}^{*}$+ CNN, and DS-RNN mix humans and obstacles in both input representations and navigation algorithm. HomoGAT only mix humans and obstacles in algorithm but not in input, while ours distinguishes them in both input and algorithm.
@@ V-B2 Ablations
V-B3 Training
V-B4 Evaluation
We test all methods with the same $500$ random unseen test episodes. For RL-based methods, we test the last 5 checkpoints (equivalent to the last 6000 training steps) and report the result of the checkpoint with the highest success rate. We conduct the testing in 5 different human and obstacle densities, as shown in the first columns of Table I and Table II. The first set of densities is the same as the training environment and is thus in training distribution. To test the generalization of methods, we change the range of human or obstacle numbers in the remaining four environments to values that do not overlap with those used in training. Thus, these 4 environments are out-of-distribution (OOD).
TABLE I: Baseline comparison results with different human and obstacle densities in unseen environments
@@ V-B4 Evaluation
TABLE II: Ablation study results with different human and obstacle densities in unseen environments
@@
Figure 6: Comparison of different methods in the same testing episode in More Constrained environment. The robot is centered in white circles and its orientation is denoted by white arrows. More qualitative results can be found in the video attachment and at https://sites.google.com/view/crowdnav-height/home.
@@
V-C Results
V-C1 Effectiveness of scene representation
To analyze the effects of input scene representations on crowd navigation algorithms, we compare ours and HomoGAT, the two methods that distinguish human and obstacle inputs, with baselines that mix humans and obstacles in input representation: the model-based DWA, the RL-based DS-RNN, and the hybrid planner $\textrm{A}^{*}$+ CNN. The results are shown in Table I.
Similarly, by representing obstacles as groups of circles, DS-RNN performs better in the More crowded environment (Fig 7(d)) than in the More constrained environment (Fig 6(d)). In addition, Table I shows that DS-RNN achieves the highest average collision rate with obstacles (0.13) in all environments. Furthermore, the obstacle collision rate of DS-RNN increases in all 4 OOD environments, indicating an overfitting problem. Among the OOD environment, the percentage of obstacle collision increase is higher in environments with higher obstacle-to-human ratios. For example, in Less crowded with the highest percentage of obstacles, the obstacle collision rate (0.21) increases by 133% w.r.t. the obstacle collision rate in training distribution (0.09). On the contrary, in More crowded with the lowest percentage of obstacles, the obstacle collision rate (0.06) drops by 33%. The reason is that DS-RNN has trouble inferring the geometric shapes of obstacles from a large group of circles, and thus fails to avoid collision with them. Thus, treating both humans and obstacles as circles is not an ideal input representation for robot navigation algorithms.
For $\textrm{A}^{*}$+ CNN, the $\textrm{A}^{*}$ global planner does not account for humans and the CNN local planner represents all observations as a 2D point cloud. As a result, from Table I, $\textrm{A}^{*}$+ CNN has the highest average timeout rate (0.10) and the longest average time (18.99) among RL-based methods in all environments, especially the two most challenging ones, for the following 2 reasons. (1) The waypoints can fail to guide the robot to reach the goal because the waypoints lose optimality as agents move. Consequently, in OOD scenarios such as Fig. 6(b), the CNN policy is especially bad at long-horizon planning because it is overfitted with good waypoints. (2) By mixing humans and obstacles in its input, the robot policy sometimes has a hard time distinguishing dynamic and static obstacles. For example, in Fig. 6(b), the robot keeps a large distance from everything and thus is less agile and efficient compared with the robots in Fig. 6(e) and (f). Thus, for RL-based approaches, treating humans as obstacles leads to overfitting to specific types of scenarios, preventing the policies from generalizing to OOD scenarios which are common in the real-world.
By splitting human and obstacle input representations, HomoGAT and HEIGHT achieve the top 2 performance across most metrics and environments. Especially, HEIGHT learns to keep a larger distance from human paths yet a shorter distance from obstacles to balance safety and efficiency (Fig. 6(f)), indicated by the lowest overall collision rates (0.17) and the shortest average navigation time (18.10) in all environments. Therefore, we conclude that our split representation consisting of detected human states and obstacle point clouds is most suitable to learn robot navigation in crowded and constrained environments with RL, among all popular choices in Table I.
V-C2 Effectiveness of the heterogeneous st-graph
To justify our heterogeneous st-graph formulation, we compare our method with $\textrm{A}^{*}$+ CNN with no graph concept and HomoGAT with a homogenous graph attention network in Table I.
$\textrm{A}^{*}$+ CNN lacks structure in input representation, and subsequently, lacks structure in network architecture. For this reason, as we discussed in Sec. V-C1, $\textrm{A}^{*}$+ CNN shows much worse average success rate (0.62 v.s. 0.84), navigation time (25.32 v.s. 18.52), and path length (12.32 v.s. 10.40) than HomoGAT which has a simple st-graph in all environments. $\textrm{A}^{*}$+ CNN also exhibits larger variances in terms of success rate (0.019 v.s. 0.0060), navigation time (3.62 v.s. 1.70), and path length (0.22 v.s. 0.019) across different OOD environments. Especially in the challenging More crowded and More constrained environments, $\textrm{A}^{*}$+ CNN exhibits a larger drop in average success rate compared with HomoGAT (0.165 v.s. 0.10). This finding shows that even a simple st-graph structure can lead to a performance gain, indicating the importance of spatial and temporal reasoning for robot crowd navigation.
However, occasional failures of HomoGAT still exist such as Fig. 6(b). Without differentiating RH and OA interactions, the robot maintains similar distances from humans and obstacles, yet fails because RH interactions require more space. As a step further, with a heterogeneous st-graph, HEIGHT treats the 3 types of edges with different network components. As a result, HEIGHT outperforms HomoGAT and achieves the best average success rate (0.87), navigation time (18.06), and path length (10.33) in all environments. The variance success rate (0.0040) and navigation time (1.40) are also the best, whereas the path length variance (0.040) is the runner-up across all environments. In addition, the average success rate drop in the two challenging environments of HEIGHT is smaller than HomoGAT (0.07 v.s. 0.10). The reason is that the heterogeneous components allow the robot to reason about the different effects of HH, RH, and OA spatial interactions. For example, in Fig. 6(f) and Fig. 7(e), HEIGHT chooses a path that avoids the most crowded region, yields to humans when the robot must encounter them, and stays closer to walls for efficiency. Therefore, we conclude that the spatial and temporal reasoning on different types of interactions is the key to ensuring good in-distribution performance and OOD generalization in crowded and interactive environments.
@@ V-C3 Importance of attention networks
We use ablations to evaluate the contribution of RH and HH attention networks to the performance of HEIGHT, as shown in Table II.
First, in terms of all metrics in most environments, the model with both RH and HH attention shows the strongest results, followed by the models with only one attention, and finally by the model with no attention. For example, in Fig. 7, the ablations in (a), (b), and (c) fail yet our full model in (e) succeeds. Especially in the 2 most challenging More crowded and More constrained environments, the existence of either RH or HH attention significantly boosts the average success rate by 0.38 and 0.43 respectively, compared with No attn. For example, in Fig. 6 (e), the robot detours and goes dangerously close to the crowds in a constrained area, yet HH attention allows it to realize the danger and resume to a less crowded path. This result shows that reasoning about both HH and RH spatial relationships is essential for our problem, especially in more crowded or more constrained environments where the spatial interactions are also dense.
@@ V-C4 Failure cases
By visualizing the testing episodes across all environments, we find that HEIGHT typically collides when (1) a human arrives at its goal and suddenly changes directions to a new goal, as shown in Fig. 8(a), or (2) the robot surroundings are extremely crowded and constrained, and almost all free paths are blocked by humans (Fig. 8(b)). In these cases, due to its speed limit, the robot sometimes cannot switch to an alternative path in time to prevent collisions. To remedy the difficulty of RL in long-term decision making [59], in future work, our method can be combined with long-term prediction and path planning algorithms that consider the stochasticity of pedestrian motions.
V-C5 Additional insights
Besides answering the four research questions above, we provide additional insights obtained from simulation experiments below.
Effectiveness of RL planning: In Table I, DWA is a model-based approach that determines robot actions based on only the current state, which leads to the worst performance. Thus, reactive policies are not sufficient to solve our problem, justifying the necessity of long-sighted planning. In contrast, $\textrm{A}^{*}$+ CNN combines planning and RL, yet relies on occupancy maps with only obstacles and not humans. As humans move, the increasingly inaccurate occupancy map reduces the optimality of waypoints, which negatively affects overall navigation. On the other hand, RL learns to maximize the expected long-term returns based on both the current and historical states of all components in the scene. Consequently, the remaining 3 RL methods, especially HomoGAT and HEIGHT, outperform $\textrm{A}^{*}$+ CNN in most metrics and in most environments. Thus, to ensure the best navigation performance, it is important to optimize the entire planning system based on the task setting instead of optimizing only a part of it.
Difficulty of scenarios: From Table I and Table II, we observe that, besides the overfitted DS-RNN, all methods show the same trend of performance change in the 5 environments: Less crowded $>$ Less constrained $>$ More constrained $>$ More crowded. Obviously, adding more humans or obstacles increases task difficulty. But interestingly, the change in the number of humans has a larger effect on the task difficulty than the change in the number of obstacles. This phenomenon shows that avoiding collisions with humans is more difficult than avoiding obstacles in nature.
Figure 9: A testing episode of our method in the real Hallway environment. The blue arrow denotes the robot path that results from its actions. The red star denotes the goal position. In this episode, the turtlebot avoids two pedestrians, one after another in a narrow corridor, enters a narrow doorway, and arrives at the goal. More qualitative results can be found in the video attachment and at https://sites.google.com/view/crowdnav-height/home.
@@ V-C5 Additional insights
TABLE III: Real-world results in 2 everyday environments
@@
VI Real-world Experiments
VI-A Experiment setup
We train the sim2real policies in the simulators of a hallway (Fig. 5(b)) and a lounge (Fig. 5(c)) for $2\times 10^{8}$ timesteps with a decaying learning rate $5\times 10^{-5}$. To learn robust policies, we inject noises into the agent positions and robot control. Then, as shown in Fig. 5 (d) and (e), we test the policies in the two corresponding everyday environments in a university building without any additional training. In the hallway environment where the free space is extremely narrow, we test the robot’s ability to handle constraints with low density crowds. In the lounge environment, we test the robot’s ability to avoid dense crowds and obstacles with more diverse shapes. The distance between the starting and the goal position of the robot ranges from 6m to 11m. The pedestrians were told to react naturally to the robot based on their own preferences. In some testing episodes, other pedestrians who were unaware of our experiment also engaged with the robot.
@@ VI-B Results
In the highly constrained yet less crowded Hallway environment, ROS navigation stack often needs to spin in place to replan, as shown by the higher navigation time in Table III. Some of the spinning recovery attempts fail and result in timeouts. In ROS navigation stack, both global and local planners treat humans as obstacles. As a result, similar to the baselines in Sec. V-C1, the robot has difficulties distinguishing dynamic obstacles that will clear out and static obstacles that will always stay, causing failures in highly constrained spaces. On the contrary, by using different representations of humans and obstacles, our method is able to explore the different strategies to avoid them through trial and error during RL training. Consequently, the robot demonstrates high success rates in Table III. Qualitatively, the robot is able to safely pass a human in a narrow corridor (Fig. 9 (a)-(d)), takes a wide turn to give a human enough room, and enters an extremely narrow doorway (Fig. 9 (e)-(h)).
In the less constrained yet more crowded Lounge environment, ROS navigation stack and ours have the same success and collision rate, because the Lounge environment has enough space for the navigation stack to apply a “stop and wait” strategy for humans and obstacles. However, as a cost, “stop and wait” takes a longer average time to arrive at goals as shown in the last column of Table III. Different from this naïve strategy, HEIGHT adapts its navigation behaviors based on different types of spatio-temporal interactions. For example, in Fig. 10 (a)-(c), since pedestrians walking in opposite directions are crossing each other, the robot first turns left to avoid the lady on its right, and then turns right to avoid the two males on its left. Then, in Fig. 10 (d)-(e) and (f)-(g), the robot chooses actions that deviate from the current and intended paths of pedestrians, walls, and the table to safely arrive at the goal. The failure cases of both methods are caused by the tables with irregular 3D shapes. Due to the height of the 2D LiDAR, some wide parts of the tables are not detected or mapped, and thus the robot cannot avoid them. Further pre-processing of the LiDAR point clouds or adding 3D sensors can mitigate this problem.

### v2

@@ V Simulation Experiments
TABLE I: Comparison of scene representation and interaction modeling of all methods
Method Scene Representation Navigation algorithm
DWA [1] Groups of small circles Treats humans as obstacles, no interaction modeling
$\textrm{A}^{*}$+ CNN [7] LiDAR point cloud A* + RL policy with no interaction modeling
DS-RNN [8] Groups of small circles RL policy with only RH attention and GRU
DRL-VO [30] Occupancy map + LiDAR point cloud Pure pursuit algorithm + RL policy with no interaction modeling
HomoGAT [45] Obstacle point cloud + detected human circles RL policy derived from st-graph with shared attention weights for all edges
Ours Obstacle point cloud + detected human circles RL policy derived from st-graph with separate functions for all edges
@@
V-A Simulation environment
We conduct simulation experiments in random environment in Fig. 5(a) developed with PyBullet [60]. The robot, humans, and static obstacles are in a 12 m×$\times$12 m arena. In each episode, rectangular obstacles are initialized with random shapes and random poses. The width and the length of each static obstacle are sampled from $\mathcal{N}(1,0.6^{2})$ and are limited in $[0.1,5]$ meters. The initial positions of the humans and the robot are also randomized. The starting and goal positions of the robot are randomly sampled inside the arena. The distance between the robot’s starting and the goal positions is between 5 m and 6 m. Some humans are dynamic and some are static. The goals of dynamic humans are set on the opposite side of their initial positions to create circle-crossing scenarios. In training, the number of humans varies from 5 to 9 and the number of static rectangular obstacles varies from 8 to 12. Among all humans, 0-2 of them are static and the rest are dynamic. In testing, the number of humans and obstacles are shown in the first column of Table II. The heights of humans and obstacles are all above the height of the robot LiDAR to ensure detectability.
@@ V-B1 Baselines
We compare the performance of our method with classical and state-of-the-art baselines listed in the top 5 row of Table I. The baselines vary in terms of both scene representation and navigation algorithm. A detailed description of all baselines can be found in Appendix-A.
@@ V-B2 Ablations
TABLE II: Baseline comparison results with different human and obstacle densities in unseen environments
@@
V-B3 Training
For fair comparision, we train all RL methods, including all baseines and ablations except DWA, with the same number of timesteps and the same PPO hyperparameters. Please refer to Appendix-B for training details.
V-B4 Evaluation
We test all methods with the same $500$ random unseen test episodes. For RL-based methods, we test the last 5 checkpoints (equivalent to the last 6000 training steps) and report the result of the checkpoint with the highest success rate. We conduct the testing in 5 different human and obstacle densities, as shown in the first columns of Table II and Table III. The first set of densities is the same as the training environment and is thus in training distribution. To test the generalization of methods, we change the range of human or obstacle numbers in the remaining four environments to values that do not overlap with those used in training. Thus, these 4 environments are out-of-distribution (OOD).
V-C Results
V-C1 Effectiveness of scene representation
To analyze the effects of input scene representations on crowd navigation algorithms, we compare ours and HomoGAT, the two methods that distinguish human and obstacle inputs, with baselines that mix humans and obstacles in input representation: the model-based DWA, the RL-based DS-RNN, and the hybrid planner $\textrm{A}^{*}$+ CNN. The results are shown in Table II.
Similarly, by representing obstacles as groups of circles, DS-RNN performs better in the More crowded environment (Fig. 7(d)) than in the More constrained environment (Fig. 6(d)). In addition, Table II shows that DS-RNN achieves the highest average collision rate with obstacles (0.13) in all environments. Furthermore, the obstacle collision rate of DS-RNN increases in all 4 OOD environments, indicating an overfitting problem. Among the OOD environment, the percentage of obstacle collision increase is higher in environments with higher obstacle-to-human ratios. For example, in Less crowded with the highest percentage of obstacles, the obstacle collision
Figure 6: Comparison of different methods in the same testing episode in More Constrained environment. The robot is centered in white circles and its orientation is denoted by white arrows. More qualitative results can be found in the video attachment.
rate (0.21) increases by 133% w.r.t. the obstacle collision rate in training distribution (0.09). On the contrary, in More crowded with the lowest percentage of obstacles, the obstacle collision rate (0.06) drops by 33%. The reason is that DS-RNN has trouble inferring the geometric shapes of obstacles from a large group of circles, and thus fails to avoid collision with them. Thus, treating both humans and obstacles as circles is not an ideal input representation for robot navigation algorithms.
For $\textrm{A}^{*}$+ CNN, the $\textrm{A}^{*}$ global planner does not account for humans and the CNN local planner represents all observations as a 2D point cloud. As a result, from Table II, $\textrm{A}^{*}$+ CNN has the highest average timeout rate (0.10) and the longest average time (18.99) among RL-based methods in all environments, especially the two most challenging ones, for the following 2 reasons. (1) The waypoints can fail to guide the robot to reach the goal because the waypoints lose optimality as agents move. Consequently, in OOD scenarios such as Fig. 6(b), the CNN policy is especially bad at long-horizon planning because it is overfitted with good waypoints. (2) By mixing humans and obstacles in its input, the robot policy sometimes has a hard time distinguishing dynamic and static obstacles. For example, in Fig. 6(b), the robot keeps a large distance from everything and thus is less agile and efficient compared with the robots in Fig. 6(e) and (f). Thus, for RL-based approaches, treating humans as obstacles leads to overfitting to specific types of scenarios, preventing the policies from generalizing to OOD scenarios which are common in the real-world.
TABLE III: Ablation study results with different human and obstacle densities in unseen environments
@@ V-C1 Effectiveness of scene representation
Compared with $\textrm{A}^{*}$+ CNN, the success rate of DRL-VO is not always better but degrades less in More crowded, Less constrained, and More constrained scenarios. This difference can be attributed to the input representation of its RL policy, which includes low-level human states encoded as occupancy maps in addition to point cloud data. Since the size of occupancy maps remains fixed regardless of the number of pedestrians, the policy generalizes well across varying human densities. However, occupancy maps have a much higher dimensionality compared to point clouds or detected human circles, which can lead to underfitting under the same training budget as HomoGAT and our method (Fig. 6(f)).
By splitting human and obstacle input representations, HomoGAT and HEIGHT achieve the top 2 performance across most metrics and environments. Especially, HEIGHT learns to keep a larger distance from human paths yet a shorter distance from obstacles to balance safety and efficiency (Fig. 6(g)), indicated by the lowest overall collision rates (0.17) and the shortest average navigation time (18.10) in all environments. Therefore, we conclude that our split representation consisting of detected human states and obstacle point clouds is most suitable to learn robot navigation in crowded and constrained environments with RL, among all popular choices in Table II.
V-C2 Effectiveness of the heterogeneous st-graph
To justify our heterogeneous st-graph, we compare our method with $\textrm{A}^{*}$+ CNN and DRL-VO with no graph concept and HomoGAT with a homogenous graph in Table II.
$\textrm{A}^{*}$+ CNN lacks structure in input representation, and subsequently, lacks structure in network architecture. For this reason, as we discussed in Sec. V-C1, $\textrm{A}^{*}$+ CNN shows much worse average success rate (0.62 v.s. 0.84), navigation time (25.32 v.s. 18.52), and path length (12.32 v.s. 10.40) than HomoGAT which has a simple st-graph in all environments. $\textrm{A}^{*}$+ CNN also exhibits larger variances in terms of success rate (0.019 v.s. 0.0060), navigation time (3.62 v.s. 1.70), and path length (0.22 v.s. 0.019) across four OOD environments. Especially in the challenging More crowded and More constrained environments, $\textrm{A}^{*}$+ CNN exhibits a larger drop in average success rate compared with HomoGAT (0.165 v.s. 0.10). A similar phenomenon can also be observed by comparing DRL-VO, which only lacks structure in network architecture, and HomoGAT. This finding shows that even a simple st-graph structure can lead to a performance gain, indicating the importance of spatial and temporal reasoning for constrained robot crowd navigation.
However, occasional failures of HomoGAT still exist such as Fig. 6(b). Without differentiating RH and OA interactions, the robot maintains similar distances from humans and obstacles, which is prone to collisions because it needs to keep larger distances from humans. As a step further, with a heterogeneous st-graph, HEIGHT treats the 3 types of edges with different network components. As a result, HEIGHT outperforms HomoGAT and achieves the best average success rate (0.87), navigation time (18.06), and path length (10.33) in all environments. HEIGHT’s variance of most metrics across all testing environments are the smallest, indicating its robustness to distribution shift. The reason is that the heterogeneous components allow the robot to reason about the different effects of HH, RH, and OA spatial interactions. For example, in Fig. 6(f) and Fig. 7(e), HEIGHT chooses a path that avoids the most crowded region, yields to humans when the robot must encounter them, and stays closer to walls for efficiency. Therefore, we conclude that the spatial and temporal reasoning on different types of interactions is the key to ensuring good in-distribution performance and OOD generalization in crowded and interactive environments.
@@ V-C3 Importance of attention networks
We use ablations to evaluate the contribution of RH and HH attention networks to the performance of HEIGHT, as shown in Table III. A visualization of attention weights can be found in Fig. 10 in Appendix.
First, in terms of all metrics in most environments, the model with both RH and HH attention shows the strongest results, followed by the models with only one attention, and finally by the model with no attention. For example, in Fig. 7, the ablations in (a), (b), and (c) fail yet our full model in (e) succeeds.Especially in the 2 most challenging More crowded and More constrained environments, the existence of either RH or HH attention significantly boosts the average success rate, compared with No attn. For example, in Fig. 6 (e), the robot detours and goes dangerously close to the crowds in a constrained area, yet HH attention allows it to realize the danger and resume to a less crowded path. This result shows that reasoning about both HH and RH spatial relationships is essential for our problem, especially in more crowded or more constrained environments where the spatial interactions are also dense.
@@ V-C4 Failure cases
By visualizing the testing episodes across all environments, we find that HEIGHT typically collides when (1) a human arrives at its goal and suddenly changes directions to a new goal (Fig. 11(a) in Appendix) or (2) the robot surroundings are extremely crowded and constrained, and almost all free paths are blocked by humans (Fig. 11(b)). In these cases, due to its speed limit, the robot sometimes cannot switch to an alternative path in time to prevent collisions. To remedy the difficulty of RL in long-term decision making [61], in future work, our method can be combined with long-term prediction and path planning algorithms that consider the stochasticity of pedestrian motions.
Besides answering the four research questions in Sec. V, we provide additional insights obtained from simulation ex- periments in Appendix-C.
Figure 8: A testing episode of our method in the real Hallway environment. The blue arrow denotes the robot path that results from its actions. The red star denotes the goal position. In this episode, the turtlebot avoids two pedestrians, one after another in a narrow corridor, enters a narrow doorway, and arrives at the goal. More qualitative results can be found in the video attachment.
@@
VI Real-world Experiments
VI-A Experiment setup
We train the sim2real policies in the simulators of a hallway (Fig. 5(b)) and a lounge (Fig. 5(c)) for $2\times 10^{8}$ timesteps with a decaying learning rate $5\times 10^{-5}$. To learn robust policies, we inject noises into the agent positions and robot control. Then, as shown in Fig. 5 (d) and (e), we test the policies in the two corresponding everyday environments in a university building without any additional training in the real-world. In the hallway environment where the free space is extremely narrow, we test the robot’s ability to handle constraints with low density crowds. In the lounge environment, we test the robot’s ability to avoid dense crowds and obstacles with more diverse shapes. Details about pedestrians and hardware configuration can be found at Appendix -D.
TABLE IV: Real-world results in 2 everyday environments
@@ VI-B Results
In the highly constrained yet less crowded Hallway environment, ROS navigation stack often needs to spin in place to replan, as shown by the higher navigation time in Table IV. Some of the spinning recovery attempts fail and result in timeouts. In ROS navigation stack, both global and local planners treat humans as obstacles. As a result, similar to the baselines in Sec. V-C1, the robot has difficulties distinguishing dynamic obstacles that will clear out and static obstacles that will always stay, causing failures in highly constrained spaces. On the contrary, by using different representations of humans and obstacles, our method is able to explore the different strategies to avoid them through trial and error during RL training. Consequently, the robot demonstrates high success rates in Table IV. Qualitatively, the robot is able to safely pass a human in a narrow corridor (Fig. 8 (a)-(d)), takes a wide turn to give a human enough room, and enters an extremely narrow doorway (Fig. 8 (e)-(h)).
In the less constrained yet more crowded Lounge environment, ROS navigation stack and ours have the same success and collision rate, because the Lounge environment has enough space for the navigation stack to apply a “stop and wait” strategy for humans and obstacles. However, as a cost, “stop and wait” takes a longer average time to arrive at goals as shown in the last column of Table IV. Different from this naïve strategy, HEIGHT adapts its navigation behaviors based on different types of spatio-temporal interactions. For example, in Fig. 9 (a)-(c), since pedestrians walking in opposite directions are crossing each other, the robot first turns left to avoid the lady on its right, and then turns right to avoid the two males on its left. Then, in Fig. 9 (d)-(e) and (f)-(g), the robot chooses actions that deviate from the current and intended paths of pedestrians, walls, and the table to safely arrive at the goal. The failure cases of both methods are caused by the tables with irregular 3D shapes. Due to the height of the 2D LiDAR, some wide parts of the tables are not detected or mapped, and thus the robot cannot avoid them. Further pre-processing of the LiDAR point clouds or adding 3D sensors can mitigate this problem.